a baseball diamond is a square with side of 90 feet. what is the shortest distanst to the nearest tenth of a foot between first and third base?
Answer by checkley79(3341)

A^2+B^2=C^2
90^2+90^2=C^2
8100+8100=C^2
16200=C^2
C=SQRT16200
C=127.3 FEET ANS.